|my.osc.edu is NOT available||Account Management||Resolved||
my.osc.edu has not been fully restored after yesterday's downtime. You can change your password, but you will not be able to use the new password on my.osc.edu. The updated password will work to... Read more
|3 days 12 hours ago||2 days 8 hours ago|
|"pbsdcp" is not working on Oakley||Oakley||Resolved||
12:35PM 5/24/2017 Update:
|3 days 11 hours ago||3 days 10 hours ago|
|Issue with GPFS on Owens since April 14, 2017||Batch, filesystem, Owens||Resolved||
3:10PM 4/18/2017 Update: Rolling reboots on Owens have started to address this GPFS issue.
We have had issues with GPFS mounts on Owens Cluster since Friday afternoon,... Read more
|1 month 1 week ago||3 weeks 4 days ago|
|Rolling reboot of all clusters, starting from Wednesday morning, April 19, 2017||Batch, Maintenance, Oakley, Owens, Ruby||Resolved||
1:40PM 4/27/2017 Update: Rolling reboots are completed.
3:10PM 4/18/2017 Update: Rolling reboots on Owens have started to address GPFS errors occured... Read more
|1 month 1 week ago||3 weeks 5 days ago|
|Scratch and Project are hung; schedulings have been paused||Batch, filesystem||Resolved||
1:00PM 4/6/2017 Update: The Scratch and Project file systems are back to normal service. Scheduling on systems are resumed. We are still investigating the causes to this problem... Read more
|1 month 3 weeks ago||1 month 3 weeks ago|
|Owens is in Partial Service||Owens||Resolved||
3:45PM April 3, 2017 Update: GPU nodes on Owens are available.
206 Owens nodes are not accessible to users due to GPU testing and a bad Ethernet switch. It is expected... Read more
|1 month 4 weeks ago||1 month 3 weeks ago|
|Rolling reboot of compute and login nodes of all clusters, starting from Wednesday morning, March 22, 2017||login, Oakley, Owens, Ruby||Resolved||
4:56PM 3/28/2017 Update: The rolling reboots of all systems are completed.
All compute nodes and login nodes of Owens, Oakley, and Ruby clusters will need to be rebooted... Read more
|2 months 1 week ago||2 months 5 hours ago|
|Update on 02/24/2017: All services available||Outage||Resolved||
02/24/17 3:50PM Update: All Services have been restored including:
|3 months 4 days ago||2 months 2 weeks ago|
|Critical change about using $PFSDIR directory at OSC||Batch||Resolved||
Starting from Thursday, Feb 2nd, the
|3 months 3 weeks ago||3 months 1 week ago|
|Issues with VDI through OnDemand||Batch||Resolved||
Update: Jan 23, 2017 3PM: this issue has been fixed.
There is an issue with OSC OnDemand -> Desktops -> Virtual Desktop Interface (VDI) such that you get "qsub: submit error ..."... Read more
|4 months 4 days ago||4 months 4 days ago|